OpenStudy (anonymous):

7/9+9/11=? fraction answer plz

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{a\times d+b\times c}{b\times d}\]

OpenStudy (isaiah.feynman):

\[\frac{ a }{ b }+\frac{ c }{ d }= \frac{ ad+bc }{ bd }\]

OpenStudy (anonymous):

can i get a staitup answer not a harder version

OpenStudy (anonymous):

\[\frac{7}{9}+\frac{9}{11}\] \[=\frac{7\times 11+9\times 9}{9\times 11}\]

OpenStudy (anonymous):

there is no royal road to arithmetic you just have to compute it

OpenStudy (anonymous):

or 77/99 + 81/99 158/99

OpenStudy (anonymous):

1 59/99

OpenStudy (anonymous):

what is so hard about saying that. damn

OpenStudy (isaiah.feynman):

@Liam4Life that formula we gave only works when the denominator of the two fractions are different.

OpenStudy (anonymous):

well still. you could of did the easy way

OpenStudy (isaiah.feynman):

That's the easiest way. :P

OpenStudy (anonymous):

multiply 9 and 11 7 and 11 9 and 9 then you have 77/99 + 81/99 then thats 1 59/99 and thats the answer

OpenStudy (anonymous):

see easy

OpenStudy (isaiah.feynman):

Easy, but not easier.